OpenStudy (dannyrod2000):

Jose is traveling up the street from his house to a local store which is 200m away. He has already started walking toward the store and is 50m away from his house. Assuming he walks at a constant velocity of 2 m/s , how long will it take to get to the store?

OpenStudy (prathamesh_m):

First find the distance from the shop. The to find the time divide that distance by his velocity.

OpenStudy (samigupta8):

Can you show me your Attempt??

OpenStudy (polymorphism):

d= rt d= (200-50) r=2 We need to find the t = (200 - 50) / 2